OpenStudy (anonymous):

HARD MATH QUESTION FOR MATH EXPERTS A girl went to a horse race and she forgot to count the horses and humans there, but she knows there are 196 legs and 74 heads. How many horses and humans were there at the horse race?

OpenStudy (anonymous):

@Directrix @mathstudent55 @Hero

OpenStudy (sw050399):

Well, you'd start with writing 2 equations that express what you know,

OpenStudy (anonymous):

It's not a question for school or anything, i just wanted to put some minds to work, can you figure it out? :)

OpenStudy (mathstudent55):

Let h = number of horses, and let p = number of people. How many legs do h horses have? How many legs do p people have?

OpenStudy (anonymous):

4 h per horse 2 per human of course

OpenStudy (mathstudent55):

h horses have 4h legs p people have 2p legs

OpenStudy (mathstudent55):

We know there are 196 legs in total, so 4h + 2p = 196

hero (hero):

Where's the "hard" question?

OpenStudy (anonymous):

there would be 24 horses and 50 humans. @sw050399 Close.. keep trying

OpenStudy (mathstudent55):

Since both horses and people have one head each, there are h + p headds in total, but we know there are 74 heads, so we have h + p = 74

hero (hero):

I hardly consider this hard

hero (hero):

I've seen far worse

OpenStudy (mathstudent55):

Now you have two equations and two variables. 4h + 2p = 196 h + p = 74 Do you know how to solve a system of equations?

mathslover (mathslover):

51 humans and 24 horses

OpenStudy (anonymous):

How'd you figure it out so fast?

mathslover (mathslover):

Girl would also be included

OpenStudy (anonymous):

Troll.

OpenStudy (anonymous):

exactly, thats where @sw050399 forgot to count. and @Hero Sorry, it took me a while to understand

mathslover (mathslover):

So is your doubt clear now? :P

OpenStudy (anonymous):

@mangnagelytis, not at all, I wanted to see if someone would understand this question as I.

mathslover (mathslover):

\(2x + 4y = 196\) \(x + y = 74\) \(x + 3y = 122\) y = 24 Thus x = 50 but +1 there as there is a girl also.

OpenStudy (anonymous):

what about this one? X 0 0 X The figure above represents four offices that will be assigned randomly to four employees, one employee per office. If Karen and Tina are two of the four employees, what is the probability that each will be assigned an office indicated with an X? a. 1/16 b. 1/12 c. 1/6 d. 1/4 e. 1/2

OpenStudy (anonymous):

(I found this on a math practice website, supposedly "extremely hard."

OpenStudy (mathstudent55):

This is how you solve the system of equations by elimination. Multiply the second equation by -2 and write it below the first equation. 4h + 2p = 196 -2h - 2p = -148 ----------------------(add the equations) 2h = 48 h = 24 (24 horses) Now substitute 24 for h in the original second equation and solve for p: h + p = 74 h + 24 = 74 h = 50 (50 people)

OpenStudy (anonymous):

@mathstudent55 51 people is correct answer.

OpenStudy (mathstudent55):

How many legs do 51 people have?

mathslover (mathslover):

I was thinking the same @mathstudent55

OpenStudy (mathstudent55):

50 people @ 2 legs per person = 100 legs 24 horses @ 4 legs per horse = 96 legs ----------------------------------- 74 heads 196 legs

OpenStudy (anonymous):

You forgot to count, the girl thats where you went wrong, but your systematic equations and work was very complete and thorough.

OpenStudy (anonymous):

she didn't include herself when saying how many heads and legs she had counted while there.

mathslover (mathslover):

@iloveyou1212 51 people + 24 horses have : 102 + 96 =198 legs..

OpenStudy (mathstudent55):

The girl is included in the total count. She is a human, so she was counted with the humans.

OpenStudy (anonymous):

Well not necessarily in the question, because technically she counted what was there, and when she made the final count, she didn't include herself.. That is what my teacher told me when we went over the answers..

OpenStudy (anonymous):

@mathstudent55 You are correct, my teacher isn't bright.

OpenStudy (anonymous):

My apologies:)

OpenStudy (mathstudent55):

np

mathslover (mathslover):

Someone took something from me. :)

mathslover (mathslover):

great work @mathstudent55